An auditorium has 20 rows of seats. The first row contains 30 seats. As you move to the rear of the auditorium, each row has 5 more seats than the previous row.
How many seats are in the 13th row? .
How many seats are in the auditorium?

Answer by Boreal(15235)

This is an arithmetic series
The first row has 30 seats, and the 13th row is 12 rows behind.
N(13)=20+(13-1)*5
N=20*60=80
Sum=(1/2) {first row + last row}* n rows
There are 20 rows, and by the above formula, row 20 has 20+(19*5)=115 seats.
sum=(1/2)*20 (20+115)=1350 seats
20+25+30+35+40+45+50+55+60+65+70+75+80+85+90+95+100+105+110+115 ;;; 20 rows
adding the ones ending in 0, and get 650
adding together the ones ending in 5=700
sum is 1350